a cashier has 21 fewer dimes than nickels, 15 more quarters than nickles that worth $12.05 in total. find how many dimes she has
Answer by richwmiller(17219)

d+21=n
q-15=n
25q+10d+5n=1205
d = 5, n = 26, q = 41